OpenStudy (deadshot):

What is the simplified form of x minus 4 over x squared minus x minus 12 • x minus 4 over x squared minus 8x plus 16 ?

Directrix (directrix):

@DeadShot Please click on the blue DRAW button below and write the posted problem by hand. Thanks.

OpenStudy (unklerhaukus):

is this your expression?\[ \frac{x - 4}{ x^2 - x - 12} \times\frac{ x - 4 }{ x^2 - 8x + 16}\]

OpenStudy (unklerhaukus):

try factorising the denominators , then look to cancel terms